1. Here are data regarding Acme Corporation’s European Division:

Average operating assets $300,000

Sales revenue $400,000

Net operating income $60,000

Minimum required rate of return 14%

Compute the European Division’s: a. Turnover ___________ b. Margin ____________ c. Return on investment ____________ d. Residual income __________________
Answer #1

Average operating assets = $300,000

Sales revenue = $400,000

Net operating income = $60,000

Minimum required rate of return = 14%

a. Turnover = Sales revenue/Average operating assets

= 400,000/300,000

= 1.33 times

b. Margin = Net operating income/Sales revenue

= 60,000/400,000

= 15%

c. Return on investment = Net operating income/Average operating assets

= 60,000/300,000

= 20%

D. Residual income = Net operating income - (Average operating assets x Minimum required rate of return)

= 60,000 - (300,000 x 14%)

= 60,000 - 42,000

= $18,000

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.